Tim bought a house for $500,000 in 2005 in California. When his daughter Mary got married in August 2014, he gave the house to Mary as a gift. The fair market value of the house was $800,000. Mary and her husband Jerry have lived in the house since then.

  1. The house is Mary and Jerry’s community property (True or False)
  2. How much is the taxable gift?

If Mary and Jerry sell the house for $900,000 in April 2019:

  1. What is the amount of capital gain? Is the capital gain long term or short term?
  2. How much is the capital gain tax assuming their income tax rate is 32%?
Answer #1

NO, False the property was originally owned by Tim and subsequently gifted to Mary and jerry, hence it is not a community property.

The gift is not taxable for Mary and Jerry as they are recipients, the gift is taxable in the hands of the donor i.e. Tim has to either pay donor gift tax on Fair value of property gifted less annual limit of $15,000 ((900,000 -15,000 = $885,000) or reduce $885,000 from the life time limit.

Capital gain arises when Mary and Jerry sell the property for $ 900,000. The capital gain is Long term capital gain.

Capital gain = Consideration - cost to original owner

= $ 900,000 - $ 400,000

= $ 500,000

The long term capital in this would be 15% . As Mary and Jerry lived in it for more than 2 years, selling it will be considered as sale of primary residence. hence, an exclusion of $ 250,000 in capital gains is available if filing is single and $500,000 if filling jointly.

Therefore, if Mary and Jerry are filing jointly then Capital gains = $500,000 - $500,000 = 0

if filing is single then = $500,000 - $ 250,000 = 250,000 * 15% = $ 30,000